A bank loaned out $14,500, part of it at the rate of 12% annual interest, and the rest at 5% annual interest. The total interest earned for both loans was $1,005.00. How much was loaned at each rate? show me explanation step by step
with solution
A bookcase is to have 4 shelves including the top as pictured below.
The width is to be 17 feet less than 4 times the height. Find the width and the height if the carpenter expects to use 22 feet of lumber to make it.


Answer by VFBundy(438)

A bank loaned out $14,500, part of it at the rate of 12% annual interest, and the rest at 5% annual interest. The total interest earned for both loans was $1,005.00. How much was loaned at each rate?

12% interest portion:
Principal = p
Rate = 0.12
Interest = 0.12p

5% interest portion:
Principal = 14500 - p
Rate = 0.05
Interest = 0.05(14500 - p) = 725 - 0.05p

(0.12p) + (725 - 0.05p) = 1005

0.07p + 725 = 1005

0.07p = 280

p = 4000

12% interest portion:
Principal = p = $4000

5% interest portion:
Principal = 14500 - p = 14500 - 4000 = $10,500